OpenStudy (anonymous):

integral of e^x+lnx^2dx

OpenStudy (anonymous):

I mean \[\int\limits_{}^{} e^{x+lnx^2}dx\]

sam (.sam.):

Integration by parts

sam (.sam.):

u=x^2 dv=e^xdx du=2xdx v=e^x \[e^x x^2-2\int\limits e^x x \, dx \]

sam (.sam.):

By parts again u=x dv=e^x du=dx v=e^x

sam (.sam.):

\[e^x x^2-2 e^x x+2\int\limits e^x \, dx \]

sam (.sam.):

Solve can simplify

sam (.sam.):

I mean and*

OpenStudy (anonymous):

how do we decide u and dv in general :(

sam (.sam.):

Its integration by parts, you must let u=x^2 NOT u=e^x because when you differentiate u=x^2, it will get smaller and smaller. but if you take u=e^x then when you differentiate it it's still the same, and you'll get infinite loop